A point source emits light of only a single frequency = 6.7 x 101​ 4​ Hz. What color is the light produced?
1
Expert's answer
2020-12-02T09:49:45-0500

Let's first find the wavelength of the light wave:


"\lambda=\dfrac{c}{f}=\dfrac{3\cdot10^8\ \dfrac{m}{s}}{6.7\cdot10^{14}\ Hz}=448\cdot10^{-9}\ m=448\ nm."


Indigo color has the range of the wavelengths between 440 and 450 nanometers. Therefore, the point source produces the light of indigo color.

Answer:

"\lambda=448\ nm," indigo.